# SWAP-less Implementation of Quantum Algorithms

### v1

@@ Abstract
We present a formalism based on tracking the flow of parity quantum information to implement algorithms on devices with limited connectivity without qubit overhead, SWAP operations or shuttling. Instead, we leverage the fact that entangling gates not only manipulate quantum states but can also be exploited to transport quantum information. We demonstrate the effectiveness of this method by applying it to the quantum Fourier transform (QFT) and the Quantum Approximate Optimization Algorithm (QAOA) with $n$ qubits. This improves upon all state-of-the-art implementations of the QFT on a linear nearest-neighbor architecture, resulting in a total circuit depth of ${5n-3}$ and requiring ${n^{2}-1}$ CNOT gates. For the QAOA, our method outperforms SWAP networks, which are currently the most efficient implementation of the QAOA on a linear architecture. We further demonstrate the potential to balance qubit count against circuit depth by implementing the QAOA on twice the number of qubits using bi-linear connectivity, which approximately halves the circuit depth.
@@ II Parity label tracking
Parity codes [28, 24, 25] encode information of multiple logical qubits onto single physical qubits in order to manipulate it locally. If the multi-qubit information is given along the same axis, e.g., $Z_{0}Z_{1}Z_{2}$ or $X_{0}X_{1}$ for Pauli-operators $X_{i}$ or $Z_{i}$, it is referred to as the parity of the qubits involved. Typically, parity codes focus on localizing parities of a single basis, since these can always be localized at the same time, which in turn also leads to the appearing stabilizers being defined in the same basis. In the following we choose to focus on $Z$-parities, where the physical qubits holding multi-qubit $Z$-information are called parity qubits, while physical qubits trivially encoding single-qubit $Z$-information are referred to as base qubits  [^1] Note that the base qubits were termed data qubits in previous literature about the parity architecture, e.g., Ref. [24]..
In this section, we present the formalism of tracking parity information [30, 31, 32, 33] and show why this is a helpful tool to design efficient quantum algorithms. To this end, we assign a label $P^{(\lambda)}_{a}$ to every physical qubit $a$ at time $\tau_{\lambda}$ and define each qubit label to be the set ${P^{(0)}_{a}=\{a\}}$ at the starting point $\tau_{0}$ of the label tracking. During a time step $\lambda$, which is defined as the time period between time $\tau_{\lambda-1}$ and time $\tau_{\lambda}$, the label of a qubit $t$ can be changed by applying the gate ${\rm{CNOT}}_{{c}\rightarrow{t}}$, where the index $c$ denotes the control qubit and $t$ the target qubit of the CNOT gate. The labels are tracked accordingly as
@@
Here and in the following, logical operators are denoted with a bar. As we will see later, this also translates to general rotations, i.e., exponentials of such operators. The labels also indicate how to apply a logical $X$-operation on qubit $j$, namely by flipping all the physical qubits that encode $\bar{Z}_{j}$ information and therefore hold index $j$ in their label. Consequently, the logical operation $\bar{X}_{j}$ can be implemented physically as
@@
Figure 1: Labeling of qubits demonstrating the effects of CNOT gates on the qubit encoding. The labels $P^{(1)}_{2}$ and $P^{(2)}_{2}$ of physical qubit $2$ respectively at time $\tau_{1}$ and $\tau_{2}$ are explicitly shown and used as example in the main text. The initial labels $P_{j}^{(0)}$ directly correspond to individual logical qubit information $\bar{Z}_{j}$ labeled $0,1$ and $2$. Furthermore, the labels change at the targets of the CNOTs according to Eq. (1). The colored lines between the qubit vertices represent logical lines, indicating how the corresponding logical X-operator is spread out onto physical operators.
@@
As an example, take the three-qubit circuit which implements all possible multi-qubit $Z$-rotations, shown in Fig. 1. Therein, let us focus on physical qubit $2$, on which the desired logical operations are $e^{i\alpha\bar{Z}_{0}\bar{Z}_{1}\bar{Z}_{2}}$ and $e^{i\beta\bar{Z}_{1}\bar{Z}_{2}}$. Applying the circuit $C_{1}=\rm{CNOT}_{1\rightarrow 2}\rm{CNOT}_{0\rightarrow 1}$ changes the label ${P_{2}^{(0)}}=\{2\}$ to ${P_{2}^{(1)}=\{0,1,2\}}$, such that the physical single-qubit rotation $e^{i\alpha Z_{2}}$ effectively applies the logical multi-qubit rotation $e^{i\alpha\bar{Z}_{0}\bar{Z}_{1}\bar{Z}_{2}}$. Note that applying the inverse sequence $C^{\dagger}_{1}$ (gray in the figure) after the rotation would restore the initial labels. This view of encoding, single-qubit rotation and decoding would correspond to a phase gadget [34] of the form
@@
holds. At time ${\tau_{\lambda}}$, the labeling generalizes the logical operator corresponding to a physical single-qubit $Z$-operator of qubit $a$ to
@@
An analog expression generalizes the logical $X$-operators as defined by the logical line $\mathcal{L}_{l}^{(\lambda)}$ (and thus the labels) as
@@
Additionally, a second set of labels could be tracked or computed from the $Z$-labels for the $X$-basis specifically, which would effectively track the labels of the inverse circuit, leading to a similar expression as Eq. (9) for the $X$-basis. The rules for tracking labels for the $X$-basis, $Q^{(\lambda)}_{a}$, follow from the fact that a Hadamard-transform of the CNOT gate effectively interchanges the control and target of this CNOT gate. Furthermore, the labels can be used to express arbitrary unitary operators $\bar{U}$, by decomposing them into
$\bar{U}=\prod_{i}\bar{U}^{X}_{Q^{i}}(\theta_{i})\bar{U}^{Z}_{P^{i}}(\phi_{i}),$ (11)
@@
with $P^{i}$ and $Q^{i}$ denoting labels for the $Z$- and $X$-basis, respectively, ${\bar{U}^{Z}_{P}(\phi)=\exp\left(i\phi\,\Pi_{j\in P}\bar{Z}_{j}\right)}$ and ${\bar{U}^{X}_{Q}(\theta)=\exp\left(i\theta\,\Pi_{j\in Q}\bar{X}_{j}\right)}$. Note that multiple subsequent labels can occur at the same time $\tau_{\lambda}$, e.g., ${P^{i}=P_{a}^{(\lambda)}}$ and ${P^{i+1}=P_{b}^{(\lambda)}}$. This decomposition is always possible, since these operators are physically implemented via $X$- and $Z$-rotations and CNOT gates [see Eqs. (9) and (10)], which together form a universal gate set.
@@
Figure 2: (a) A unit cell of the extended LHZ-code for $5$ logical qubits in two spatial dimensions. The large arrows indicate the periodic boundary conditions, i.e., joining the edges indicated by the arrows forms a closed strip with periodic labeling. For example, the qubits with labels $\{0,3\}$, $\{0,4\}$, $\{1,3\}$ and $\{1,4\}$ at the boundaries form a $Z$-stabilizer. Two different spanning lines are highlighted by the qubits and lines with thick borders, reaching from the top boundary to the bottom boundary. In the spatial picture, the colored parts correspond to overlap of a logical line with a spanning line, while in the temporal encoding they correspond to the full logical lines. The horizontal arrows indicate how the spanning line effectively evolves from left to right by applying the quantum circuit in (b), which changes the qubit labels. (b) The quantum circuit that changes the parity labels in the left spanning line to the labels in the right spanning line in (a). (c) Pictorial representation of the action of the gates $C^{\prime}_{1}$ and $C^{\prime}_{2}$ on the spanning lines.
@@
Although the example in Fig. 1 does not have any stabilizers, i.e., it encodes $n$ qubits onto $n$ qubits, it can still be considered a (parity) code. In order to allow for parallelization or error correction capabilities, redundancy of the encoded information is required, which can be achieved with the help of auxiliary qubits. If these are initialized in the $+1$ $Z$-eigenstate $|0\rangle$, applying a physical $Z$-gate to these qubits has a trivial effect, as $Z|0\rangle=I|0\rangle$. In our formalism, this is represented by giving such qubits an empty set as a label and considering the empty product as an identity, such that the corresponding logical operator is also the identity, i.e., ${\Pi_{j\in\{\}}\bar{Z}_{j}=I}$. Accordingly, we call these qubits empty qubits, as opposed to active qubits, which have a non-empty label. An empty qubit can become an active qubit by targeting it with CNOT gates which are controlled by active qubits, such that a non-empty parity label is generated on the previously empty qubit. Any parity code encoding $n$ logical qubits can be built up from $n$ initial active qubits and any number of additional empty qubits, with a corresponding encoding circuit [26]. Each added empty qubit adds one stabilizer operator to the code definition. At initialization of the empty qubit, the stabilizer is the $Z$-operator of that qubit. With the encoding circuit, such single-qubit stabilizers are mapped onto multi-qubit operators, which are represented as gray polygons, as for example shown in Fig. 2(a). Note that the tracking method only remains valid as long as no physical operations that violate any stabilizers are performed.
III Quantum circuits inspired by the LHZ-code
@@
So far, we have introduced parity labels and their tracking. In the following sections, we illustrate the use of the tracking in quantum circuit design based on the LHZ-code [23, 24], which contains the labels that correspond to the interactions required for the execution of promising quantum algorithms. The LHZ-code emerges when encoding $n$ logical qubits onto additional empty qubits, such that all possible two-body parities appear, by means of tracking the labels throughout a specific encoding circuit [24]. This results in $n$ base qubits and $n(n-1)/2$ parity qubits holding all possible single- and two-body parities in their labels. Changing an empty qubit into a parity qubit induces a new stabilizer on multiple physical qubits, as described in Sec. II, which are geometrically local on a 2D-grid. The resulting LHZ-code for ${n=5}$ qubits appears twice in Fig. 2(a), shown as qubits connected by dark-gray polygons. Both sets of dark-gray polygons correspond to $n(n-1)/2$ stabilizers, which constrain the logical state space.
III.1 Spatial extension of the LHZ-code
@@
The LHZ-code can be extended in space by positioning two LHZ-codes next to each other, where one is turned upside down, such that the light gray stabilizers [see Fig. 2(a)] fuse the two LHZ-codes. This process can be repeated to create a periodic code, with a period of ${n+1}$ qubits in the horizontal direction, as indicated by the vertical arrows at the left and right boundary of Fig. 2(a). An essential property of the extended layout is that any line of physical qubits connecting the top boundary to the bottom boundary contains all logical information, i.e., the information required to construct the complete logical state vector. In the following, we will denote such sets of qubits as spanning lines. Due to the periodicity of the labels, each spanning line reoccurs after a period of ${n+1}$ qubits. However, after half a period, i.e., after at most $\lceil\frac{n+1}{2}\rceil$ qubits, every spanning line also appears mirrored around the horizontal axis located exactly in the middle between the top and bottom boundary. Therefore, the region within half a period in the horizontal direction contains all single-qubit and two-qubit $Z$ parities in its labels. This essential feature can be utilized to design a quantum circuit that applies all of these parities, but without the need for any stabilizers.
@@ III.2 Temporal parity quantum computing
At the start of the label tracking the qubits in the chain form a spanning line composed solely of base qubits. Applying a sequence of CNOT gates allows us to change the labels in order to obtain any of the other $n$-qubit spanning lines present in the extended LHZ-code. Let us consider the spanning lines highlighted by the black borders in Fig. 2(a). Starting from the labels in the spanning line on the left, the CNOT circuit in Fig. 2(b) changes the labels to the spanning line on the right. As an example, the circuit elements $C^{\prime}_{1}$ and $C^{\prime}_{2}$ change the label $\{0,4\}$ into $\{1\}$ and label $\{0\}$ into $\{0,1\}$, respectively. When considering this label change in the spatial picture of Fig. 2(a), the spanning line effectively moves across the stabilizers from left to the right, as indicated by the horizontal arrows. This picture enables the extended LHZ-code to be used as a guide to find the CNOT circuits to change from one spanning line to the other, by expressing the movement across the stabilizers as a CNOT circuit. If a spanning line contains all of the labels in a given stabilizer except for the label assigned to qubit $a$, a CNOT circuit can be used to deform the spanning line along this stabilizer, i.e., changing the label of one of the overlapping labels into the label of qubit $a$. This process is illustrated for the spatial picture in Fig. 2(c) for weight-4 and weight-3 stabilizers. Since the extended LHZ-code exclusively contains such stabilizers, this operation always results in applying one CNOT gate on the boundary or two CNOTs in the bulk, respectively. The CNOT circuit to change all the labels in a spanning line is composed of two rounds of ${n-1}$ commuting CNOT gates with at most $2$ CNOT targets or controls on the same qubit, as can be seen in Fig. 2(b). We assume that a single qubit can only be part of one gate at a time. For this reason, such a round of CNOT gates is considered to have a circuit depth of $2$.
Using the extended LHZ-code as a guide for label tracking is particularly beneficial for the implementation of algorithms that require the interactions specific to the Hamiltonian of a transverse-field Ising model, i.e., two-body $Z$-interactions and single-body $X$-interactions, for which it allows for improved circuit depth and gate count compared to other currently known methods. The upcoming section introduces explicitly how this is implemented for the QFT and the QAOA.

### v2

@@ Abstract
We present a formalism based on tracking the flow of parity quantum information to implement algorithms on devices with limited connectivity without qubit overhead, SWAP operations or shuttling. Instead, we leverage the fact that entangling gates not only manipulate quantum states but can also be exploited to transport quantum information. We demonstrate the effectiveness of this method by applying it to the quantum Fourier transform (QFT) and the Quantum Approximate Optimization Algorithm (QAOA) with $n$ qubits. This improves upon all state-of-the-art implementations of the QFT on a linear nearest-neighbor architecture, resulting in a total circuit depth of 5n3{5n-3} and requiring $n^{2}-1$} CNOT gates. For the QAOA, our method outperforms SWAP networks, which are currently the most efficient implementation of the QAOA on a linear architecture. We further demonstrate the potential to balance qubit count against circuit depth by implementing the QAOA on twice the number of qubits using bi-linear connectivity, which approximately halves the circuit depth.
@@ II Parity label tracking
Parity codes [28, 24, 25, 29] encode information of multiple logical qubits onto single physical qubits in order to manipulate it locally. If the multi-qubit information is given along the same axis, e.g., $Z_{0}Z_{1}Z_{2}$ or $X_{0}X_{1}$ for Pauli operators $X_{i}$ or $Z_{i}$, it is referred to as the parity of the qubits involved. Typically, parity codes focus on localizing parities of a single basis, since these can always be localized at the same time, which in turn also leads to the appearing stabilizers being defined in the same basis. In the following we choose to focus on $Z$ parities, where the physical qubits holding multi-qubit $Z$ information are called parity qubits, while physical qubits trivially encoding single-qubit $Z$ information are referred to as base qubits  [^1] Note that the base qubits were termed data qubits in previous literature about the parity architecture, e.g., Ref. [24]..
In this section, we present the formalism of tracking parity information and show why this is a helpful tool to design efficient quantum algorithms. The existing methods to design quantum algorithms that utilize parity information [31, 32, 33] are typically used in a heuristic manner for a given pair of algorithm and connectivity. In contrast, we provide quantum circuits for QFT and QAOA on linear architectures for any system size, leading to analytical resource estimates. To this end, we assign a label $P^{(\lambda)}_{a}$ to every physical qubit $a$ at time $\tau_{\lambda}$ and define each qubit label to be the set $P^{(0)}_{a}=\{a\}$} at the starting point $\tau_{0}$ of the label tracking. During a time step $\lambda$, which is defined as the time period between time $\tau_{\lambda-1}$ and time $\tau_{\lambda}$, the label of a qubit $t$ can be changed by applying the gate ${\rm{CNOT}}_{{c}\rightarrow{t}}$, where the index $c$ denotes the control qubit and $t$ the target qubit of the CNOT gate. The labels are tracked accordingly as
@@
Here and in the following, logical operators are denoted with a bar. As we will see later, this also translates to general rotations, i.e., exponentials of such operators. The labels also indicate how to apply a logical $X$ operation on qubit $j$, namely by flipping all the physical qubits that encode $\bar{Z}_{j}$ information and therefore hold index $j$ in their label. Consequently, the logical operation $\bar{X}_{j}$ can be implemented physically as
@@
Figure 1: Labeling of qubits demonstrating the effects of CNOT gates on the qubit encoding. The labels $P^{(1)}_{2}$ and $P^{(2)}_{2}$ of physical qubit 22 respectively at time $\tau_{1}$ and $\tau_{2}$ are explicitly shown and used as example in the main text. The initial labels $P_{j}^{(0)}$ directly correspond to individual logical qubit information $\bar{Z}_{j}$ labeled $0,1$ and $2$. Furthermore, the labels change at the targets of the CNOTs according to Eq. (1). The colored lines between the qubit vertices represent logical lines, indicating how the corresponding logical $X$ operator is spread out onto physical operators.
@@
As an example, take the three-qubit circuit which implements all possible multi-qubit $Z$ rotations, shown in Fig. 1. Therein, let us focus on physical qubit $2$, on which the desired logical operations are $e^{i\alpha\bar{Z}_{0}\bar{Z}_{1}\bar{Z}_{2}}$ and $e^{i\beta\bar{Z}_{1}\bar{Z}_{2}}$. Applying the circuit $C_{1}=\rm{CNOT}_{1\rightarrow 2}\rm{CNOT}_{0\rightarrow 1}$ changes the label ${P_{2}^{(0)}}=\{2\}$ to $P_{2}^{(1)}=\{0,1,2\}$}, such that the physical single-qubit rotation $e^{i\alpha Z_{2}}$ effectively applies the logical multi-qubit rotation $e^{i\alpha\bar{Z}_{0}\bar{Z}_{1}\bar{Z}_{2}}$. Note that applying the inverse sequence $C^{\dagger}_{1}$ (gray in the figure) after the rotation would restore the initial labels. This view of encoding, single-qubit rotation and decoding would correspond to a phase gadget [34] of the form
@@
holds. At time $\tau_{\lambda}$}, the labeling generalizes the logical operator corresponding to a physical single-qubit $Z$ operator of qubit $a$ to
@@
An analog expression generalizes the logical $X$ operators as defined by the logical line $\mathcal{L}_{l}^{(\lambda)}$ (and thus the labels) as
@@
Additionally, a second set of labels could be tracked or computed from the $Z$ labels for the $X$ basis specifically, which would effectively track the labels of the inverse circuit, leading to a similar expression as Eq. (9) for the $X$ basis. See Ref. [35] for a generalized formalism of parity labels for any Clifford circuit. The rules for tracking labels for the $X$ basis, $Q^{(\lambda)}_{a}$, follow from the fact that a Hadamard-transform of the CNOT gate effectively interchanges the control and target of this CNOT gate. Furthermore, the labels can be used to express arbitrary unitary operators $\bar{U}$, by decomposing them into
$\bar{U}=\prod_{\lambda}\prod_{j}\bar{U}^{X}_{\lambda,j}(\theta_{j})\bar{U}^{Z}_{\lambda,j}(\phi_{j}),$ (11)
@@
with the unitaries $\bar{U}^{Z}_{\lambda,j}(\phi)=\exp(i\phi\,\Pi_{k\in P^{(\lambda)}_{j}}\bar{Z}_{k})$} and $\bar{U}^{X}_{\lambda,j}(\theta)=\exp(i\theta\,\Pi_{k\in Q^{(\lambda)}_{j}}\bar{X}_{k})$}. This decomposition is always possible, since these operators are physically implemented via $X$ and $Z$ rotations and CNOT gates [see Eqs. (9) and (10)], which together form a universal gate set.
@@
Figure 2: (a) A unit cell of the extended LHZ code for 55 logical qubits in two spatial dimensions. The large arrows indicate the periodic boundary conditions, i.e., joining the edges indicated by the arrows forms a closed strip with periodic labeling. For example, the qubits with labels $\{0,3\}$, $\{0,4\}$, $\{1,3\}$ and $\{1,4\}$ at the boundaries form a $Z$ stabilizer. Two different spanning lines are highlighted by the qubits and lines with thick borders, reaching from the top boundary to the bottom boundary. In the spatial picture, the colored parts correspond to overlap of a logical line with a spanning line, while in the temporal encoding they correspond to the full logical lines. The horizontal arrows indicate how the spanning line effectively evolves from left to right by applying the quantum circuit in (b), which changes the qubit labels. (b) The quantum circuit that changes the parity labels in the left spanning line to the labels in the right spanning line in (a). (c) Pictorial representation of the action of the gates $C^{\prime}_{1}$ and $C^{\prime}_{2}$ on the spanning lines.
@@
Although the example in Fig. 1 does not have any stabilizers, i.e., it encodes $n$ qubits onto $n$ qubits, it can still be considered a (parity) code. In order to allow for parallelization or error correction capabilities, redundancy of the encoded information is required, which can be achieved with the help of auxiliary qubits. If these are initialized in the +1+1 $Z$ eigenstate $|0\rangle$, applying a physical $Z$ gate to these qubits has a trivial effect, as $Z|0\rangle=I|0\rangle$. In our formalism, this is represented by giving such qubits an empty set as a label and considering the empty product as an identity, such that the corresponding logical operator is also the identity, i.e., $\Pi_{j\in\{\}}\bar{Z}_{j}=I$}. Accordingly, we call these qubits empty qubits, as opposed to active qubits, which have a non-empty label. An empty qubit can become an active qubit by targeting it with CNOT gates which are controlled by active qubits, such that a non-empty parity label is generated on the previously empty qubit. Any parity code encoding $n$ logical qubits can be built up from $n$ initial active qubits and any number of additional empty qubits, with a corresponding encoding circuit [26]. Each added empty qubit adds one stabilizer operator to the code definition. At initialization of the empty qubit, the stabilizer is the $Z$ operator of that qubit. With the encoding circuit, such single-qubit stabilizers are mapped onto multi-qubit operators, which are represented as gray polygons, as for example shown in Fig. 2(a). Note that the tracking method only remains valid as long as no physical operations that violate any stabilizers are performed.
III Quantum circuits inspired by the LHZ code
@@
So far, we have introduced parity labels and their tracking. In the following sections, we illustrate the use of the tracking in quantum circuit design based on the LHZ code [23, 24], which contains the labels that correspond to the interactions required for the execution of promising quantum algorithms. The LHZ code emerges when encoding $n$ logical qubits onto additional empty qubits, such that all possible two-body parities appear, by means of tracking the labels throughout a specific encoding circuit [24]. This results in $n$ base qubits and n(n1)/2n(n-1)/2 parity qubits holding all possible single- and two-body parities in their labels. Changing an empty qubit into a parity qubit induces a new stabilizer on multiple physical qubits, as described in Sec. II, which are geometrically local on a 2D grid. The resulting LHZ code for n=5{n=5} qubits appears twice in Fig. 2(a), shown as qubits connected by dark-gray polygons. Both sets of dark-gray polygons correspond to n(n1)/2n(n-1)/2 stabilizers, which constrain the logical state space.
III.1 Spatial extension of the LHZ code
@@
The LHZ code can be extended in space by positioning two LHZ codes next to each other, where one is turned upside down, such that the light gray stabilizers [see Fig. 2(a)] fuse the two LHZ codes. This process can be repeated to create a periodic code, with a period of n+1{n+1} qubits in the horizontal direction, as indicated by the vertical arrows at the left and right boundary of Fig. 2(a). An essential property of the extended layout is that any line of physical qubits connecting the top boundary to the bottom boundary contains all logical information, i.e., the information required to construct the complete logical state vector. In the following, we will denote such sets of qubits as spanning lines. Due to the periodicity of the labels, each spanning line reoccurs after a period of n+1{n+1} qubits. However, after half a period, i.e., after at most $\lceil\frac{n+1}{2}\rceil$ qubits, every spanning line also appears mirrored around the horizontal axis located exactly in the middle between the top and bottom boundary. Therefore, the region within half a period in the horizontal direction contains all single-qubit and two-qubit $Z$ parities in its labels. This essential feature can be utilized to design a quantum circuit that applies all of these parities, but without the need for any stabilizers.
@@ III.2 Temporal parity quantum computing
At the start of the label tracking the qubits in the chain form a spanning line composed solely of base qubits. Applying a sequence of CNOT gates allows us to change the labels in order to obtain any of the other $n$-qubit spanning lines present in the extended LHZ code. Let us consider the spanning lines highlighted by the black borders in Fig. 2(a). Starting from the labels in the spanning line on the left, the CNOT circuit in Fig. 2(b) changes the labels to the spanning line on the right. As an example, the circuit elements $C^{\prime}_{1}$ and $C^{\prime}_{2}$ change the label $\{0,4\}$ into $\{1\}$ and label $\{0\}$ into $\{0,1\}$, respectively. When considering this label change in the spatial picture of Fig. 2(a), the spanning line effectively moves across the stabilizers from left to the right, as indicated by the horizontal arrows. This picture enables the extended LHZ code to be used as a guide to find the CNOT circuits to change from one spanning line to the other, by expressing the movement across the stabilizers as a CNOT circuit. If a spanning line contains all of the labels in a given stabilizer except for the label assigned to qubit $a$, a CNOT circuit can be used to deform the spanning line along this stabilizer, i.e., changing the label of one of the overlapping labels into the label of qubit $a$. This process is illustrated for the spatial picture in Fig. 2(c) for weight-4 and weight-3 stabilizers. Since the extended LHZ code exclusively contains such stabilizers, this operation always results in applying one CNOT gate on the boundary or two CNOTs in the bulk, respectively. The CNOT circuit to change all the labels in a spanning line is composed of two rounds of n1{n-1} commuting CNOT gates with at most $2$ CNOT targets or controls on the same qubit, as can be seen in Fig. 2(b). We assume that a single qubit can only be part of one gate at a time. For this reason, such a round of CNOT gates is considered to have a circuit depth of $2$.
Using the extended LHZ code as a guide for label tracking is particularly beneficial for the implementation of algorithms that require the interactions specific to the Hamiltonian of a transverse-field Ising model, i.e., two-body $Z$ interactions and single-body $X$ interactions, for which it allows for improved circuit depth and gate count compared to other currently known methods. The upcoming section introduces explicitly how this is implemented for the QFT and the QAOA.